after five year, the age of a father will be thrice the age of his son, whereas five years ago, he was seven times old as his son was. What is father's present age?
Answer by JulietG(1812)

F + 5 = 3(S+5)
F - 5 = 7(S-5)
.
The easiest way to do this problem is to isolate S. We'll do that by finding the value of F and substituting.
F + 5 = 3(S+5)
Subtract 5 from each side.
F = 3S + 15 - 5
F = 3S + 10
Now take that value of F and put it into the second equation.
F - 5 = 7(S-5)
(3S+10)-5=7S-35
3S+5 = 7S - 35
Subtract 3S from each side
5 = 4S - 35
Add 35 to each side
40 = 4S
Divide each side by 4
10 = S
.
If the son is 10, then in five years, the son will be 15. The father will then be 3 times that, or 45. If he is 45 in 5 years, then he is 40 now.
Let's check by working the second half.
"five years ago, he was seven times old as his son was. "
Five years ago, the son was 5. Five years ago, the father was 35.
That is indeed seven times the age.
Success!